RGB2YCBCR(1) RGB2YCBCR(1) NAME rgb2ycbcr−convertnon-YCbCrTIFFimagestoaYCbCrTIFFimage SYNOPSIS rgb2ycbcr[options]src1.tifsrc2.tif...dst.tif DESCRIPTION rgb2ycbcrconvertsRGBcolor,greyscale,orbi-levelTIFFimagestoYCbCrimagesbytransforming andsamplingpixeldata.Ifmultiplefilesarespecifiedonthecommandlineeachsourcefileis convertedtoaseparatedirectoryinthedestinationfile. Bydefault,chrominancesamplesarecreatedbysampling2by2blocksofluminancevalues;this canbechangedwiththe−hand−voptions.OutputdataarecompressedwiththeLZWcompressionscheme,bydefault;analternateschemecanbeselectedwiththe−coption.Bydefault, outputdataarecompressedinstripswiththenumberofrowsineachstripselectedsothatthe sizeofastripisnevermorethan8kilobytes;the−roptioncanbeusedtoexplicitlysetthe numberofrowsperstrip. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm,-cjpegfortheJPEGcom-pressionalgorithm,and−clzwforLempel-Ziv&Welch(thedefault). −hSetthehorizontalsamplingdimensiontooneof:1,2(default),or4. −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. −vSettheverticalsamplingdimensiontooneof:1,2(default),or4. SEEALSO tiffinfo(1),tiffcp(1),libtiff(3) October15,1995 1